# Why $w\neq-1$? Anthropic Selection in a $\Lambda$ + Axion Dark Energy Model

### v2

@@ III Discussion and Conclusions
We have found that, within the anthropically allowed region, the axion mass typically lies around ${\cal O}(10)H_{0}$ for a decay constant near the Planck scale, and it slightly decreases for smaller decay constants (see Appendix A). While the most likely value of the dark energy equation-of-state parameter $w_{0}$ lies close to −$-1$, the deviation by ${\cal O}(0.1)$ is statistically significant. This provides a natural anthropic explanation for why $w_{0}\neq-1$ may be expected, and is intriguingly consistent with recent DESI results hinting at time-varying dark energy Adame et al. (2025); Abdul Karim et al. (2025a, b). Our results thus offer a concrete realization of how the string theory landscape, together with a light axion, can shape the nature of dark energy in our universe.
@@ Acknowledgments
We thank Wen Yin for helpful discussion on the DESI results. This work is supported by JSPS Core-to-Core Program (grant number: JPJSCCA20200002) (F.T.), JSPS KAKENHI Grant Numbers 20H01894 (F.T.), 20H05851 (F.T.), 23KJ0088 (K.M.), 24K17039 (K.M.), and 25H02165 (F.T.). This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
@@ A.2 Smaller $f$
Next, we consider smaller decay constants. We show the allowed parameter regions for $f=0.1M_{\mathrm{Pl}}$ with $m=10H_{0}$, $7H_{0}$, and $5H_{0}$ in Fig. 6. For $m=5H_{0}$, $\rho_{\Lambda}$ should be near zero to realize the accelerated universe with the limited energy budget of the axion, $\rho_{\phi}\leq 2m^{2}f^{2}$. Similarly to the bottom panel in Fig. 1, the allowed value of $w_{0}$ does not deviate largely from −$-1$. For $m=7H_{0}$ and $10H_{0}$, larger $w_{0}$ is allowed for large $\rho_{\Lambda}$.
@@
Finally, we show the probability distribution for $\rho_{\mathrm{DE,i}}$ with $f=0.1M_{\mathrm{Pl}}$ and $0.2M_{\mathrm{Pl}}$. As mentioned above, the required tuning for $\theta_{\mathrm{i}}$ becomes more severe for smaller $f$. As a result, $\rho_{\mathrm{DE,i}}$ becomes the order of the upper bound $\rho_{\Lambda}^{\mathrm{max}}$ only with a negligibly small probability. Rather, $\rho_{\mathrm{DE,i}}$ typically lies around the observed value $\rho_{\mathrm{DE,obs}}$, which is the minimum order to realize the accelerated expansion. While the probability distribution is concentrated below $\rho_{\rm DE,i}=\rho_{\rm DE,obs}$ for $f=0.1M_{\mathrm{Pl}}$, it spreads around $\rho_{\rm DE,i}\sim\rho_{\rm DE,obs}$ for $f=0.2M_{\mathrm{Pl}}$.

### v3

@@ III Discussion and Conclusions
We have found that, within the anthropically allowed region, the axion mass typically lies around ${\cal O}(10)H_{0}$ for a decay constant near the Planck scale, and it slightly decreases for smaller decay constants (see Appendix A). While the most likely value of the dark energy equation-of-state parameter $w_{0}$ lies close to 1-1, the deviation by ${\cal O}(0.1)$ is statistically significant. This provides a natural anthropic explanation for why $w_{0}\neq-1$ may be expected, and is intriguingly consistent with recent DESI results hinting at time-varying dark energy Adame et al. (2025); Abdul Karim et al. (2025a, b). Our results thus offer a concrete realization of how the string theory landscape, together with a light axion, can shape the nature of dark energy in our universe.
@@ Acknowledgments
We thank Wen Yin for helpful discussion on the DESI results. We also thank the anonymous referee for suggesting a comparison with the latest observational data. This work is supported by JSPS Core-to-Core Program (grant number: JPJSCCA20200002) (F.T.), JSPS KAKENHI Grant Numbers 20H01894 (F.T.), 20H05851 (F.T.), 23KJ0088 (K.M.), 24K17039 (K.M.), and 25H02165 (F.T.). This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
@@ A.2 Smaller $f$
Next, we consider smaller decay constants. We show the allowed parameter regions for $f=0.1M_{\mathrm{Pl}}$ with $m=10H_{0}$, $7H_{0}$, and $5H_{0}$ in Fig. 6. For $m=5H_{0}$, $\rho_{\Lambda}$ should be near zero to realize the accelerated universe with the limited energy budget of the axion, $\rho_{\phi}\leq 2m^{2}f^{2}$. Similarly to the bottom panel in Fig. 1, the allowed value of $w_{0}$ does not deviate largely from 1-1. For $m=7H_{0}$ and $10H_{0}$, larger $w_{0}$ is allowed for large $\rho_{\Lambda}$.
@@
Finally, we show the probability distribution for $\rho_{\mathrm{DE,i}}$ with $f=0.1M_{\mathrm{Pl}}$ and $0.2M_{\mathrm{Pl}}$ in Fig. 8. As mentioned above, the required tuning for $\theta_{\mathrm{i}}$ becomes more severe for smaller $f$. As a result, $\rho_{\mathrm{DE,i}}$ becomes the order of the upper bound $\rho_{\Lambda}^{\mathrm{max}}$ only with a negligibly small probability. Rather, $\rho_{\mathrm{DE,i}}$ typically lies around the observed value $\rho_{\mathrm{DE,obs}}$, which is the minimum order to realize the accelerated expansion. While the probability distribution is concentrated below $\rho_{\rm DE,i}=\rho_{\rm DE,obs}$ for $f=0.1M_{\mathrm{Pl}}$, it spreads around $\rho_{\rm DE,i}\sim\rho_{\rm DE,obs}$ for $f=0.2M_{\mathrm{Pl}}$.
@@
Appendix B Comparison with observational data
Here, we compare our results with the latest observational analysis Luu et al. (2025b), which appeared on the arXiv after our submission. The works of Refs. Luu et al. (2025a, b) consider a dark energy model that consists of of a cosmological constant and an axion and obtain the best-fit values of the model parameters from cosmological data including baryon acoustic oscillations (BAO), CMB, and Type Ia supernovae. Focusing on a negative cosmological constant and an axion decay constant of $f=M_{\mathrm{Pl}}$, they find the best-fit values Luu et al. (2025b)
$\begin{gathered}\rho_{\Lambda}=-2.2\rho_{\mathrm{DE,obs}}\ ,\quad m=2.0H_{0}\ ,\\ \theta_{i}=0.18\pi\ ,\quad w_{0}=-0.74\ ,\end{gathered}$ (19)
in our notation. Since these values are obtained from observational data, the total dark energy density matches the observed value, which lies at the edge of the probability distribution in our anthropic discussion (Fig. 4). Correspondingly, the preferred axion mass also lies on the low-mass side of the peak for the likelihood of $\ln m$ shown in Fig. 2. The best-fit parameters are allowed by anthropic considerations, but they are not necessarily favored. For instance, in the case of the cosmological constant, the anthropic upper bound is about three orders of magnitude larger than the observed value. This may suggest that the anthropic conditions we have considered are not sufficient and that additional, more stringent conditions may exist, or it may simply be a coincidence. We note that, as discussed in the previous section, if the decay constant is slightly smaller than $M_{\rm Pl}$, the likelihood of the best-fit parameters can be enhanced.
We also compare our results with BAO. The observation of BAO is typically quantified by the ratio of the comoving distance $D_{M}(z)$ to the redshift $z$ and the sound horizon $r_{d}$. While $r_{d}$ is determined by the cosmic expansion before the recombination and is insensitive to the properties of dark energy, $D_{M}(z)$ is directly affected by the late-time expansion history. When we take $\rho_{\mathrm{DE,i}}=\mathcal{O}(10^{3})\rho_{\mathrm{DE,obs}}$ as a typical value in our anthropic discussion, the resultant value of $D_{M}(z)/r_{d}$ for fixed $z$ becomes much smaller than the value suggested by the observations, reflecting the larger Hubble parameter in the recent universe.
As another quantity representing the cosmological evolution, we can evaluate the lifetime of the universe in the presence of a negative cosmological constant. Using the best-fit parameters (19), the total lifetime of the universe is estimated as $T\simeq 33.3$ Gyr Luu et al. (2025b). Here, we roughly estimate the lifetime of the universe for typical parameters in our anthropic discussion. First, according to Eq. (13), the axion starts to oscillate at $t=t_{\mathrm{osc}}\sim m^{-1}$ up to logarithmic corrections. Once the axion starts to oscillate and comes to behave as non-relativistic matter, the universe recollapses in a timescale of $\mathcal{O}(H_{\mathrm{osc}}^{-1})$ with $H_{\mathrm{osc}}$ being the Hubble parameter at $t=t_{\mathrm{osc}}$ Luu et al. (2025b). Thus, the lifetime of the universe is roughly estimated as
$T\sim m^{-1}+\sqrt{\frac{\rho_{\mathrm{DE,obs}}}{\rho_{\mathrm{DE,i}}}}H_{0}^{-1}\ ,$ (20)
which is shorter than the age of the universe $\sim H_{0}^{-1}$ for typical parameters, $m=\mathcal{O}(10\,\text{--}\,100)H_{0}$ and $\rho_{\mathrm{DE,i}}=\mathcal{O}(10^{3})\rho_{\mathrm{DE,obs}}$. This implies that the universe often recollapses too early in our anthropic argument. Perhaps the emergence of intellectual life like ours requires a significantly longer cosmic lifetime, which may point to additional anthropic constraints beyond those considered here, or to a slightly smaller axion decay constant.